Hi,
Can I please have an answer for this equation
3t - 12= 4- t
B. Prove the solution is correct
Could I please have the working out shown to me as I need to understand the answer and that really helps.

Answer by Alan3354(69443)

3t - 12= 4- t
Add t
4t - 12 = 4
Add 12
4t = 16
t = 4